# On the solvable-unsolvable transition due to noise-induced chaos in digital memcomputing

### v1

@@ 2.3 Physical noise
To make contact with experimental realizations of these machines, we introduce physical noise through stochastic terms in the ODEs of DMMs. In view of the fact that in possible electrical circuit realizations of DMMs, the memory variables are the ones most susceptible to noise [1], we add stochastic terms to Eqs. (2) and (3). Of course, such noise propagates indirectly to the fast degrees of freedom $v_{n}$’s through their coupling in Eq. (1). Indeed, as discussed in [36], noise can for instance serve as a proxy for temperature, since temperature fluctuations can induce physical instabilities in the resistive memory elements of the gates of DMMs.
@@
If we indicate with $\tilde{x}_{l,m}$ the noisy version of the long-term memory variables, the stochastic terms are linked to the noiseless dynamical system’s memory by
$\dot{\tilde{x}}_{l,m}(t)=\dot{x}_{l,m}(t)+\Gamma\eta(t)$ (9)
@@
with $\eta\sim\mathcal{N}(0,1)$, a Gaussian distribution with zero average and unit variance, and $\Gamma$ is the strength of the noise. Similarly for the short term memory $x_{s,m}$. Gaussian noise is then added randomly to the memory variables at every time step with an intensity that depends on the noise strength $\Gamma$.
3 Results
3.1 Numerical noise
In numerically integrating Eqs. (1)-(3) using the forward Euler method, we transform the continuous-time dynamics into a discrete map. We view related errors induced by such a transformation as the primary components of numerical noise, which can sometimes force the system into chaotic behavior.
@@ 3.1 Numerical noise
Figure 1: Percentage of solved instances and EAMLLE ($\langle\overline{\lambda}\rangle$) vs. integration time step, $\Delta t$. Each point used to make this plot was obtained using an ensemble of 100 solvable 3-SAT instances with clause-to-variable ratio $M/N=7$ with $N=1000$. The upper bound for each trajectory was set to 20,000 steps.
@@
It is clear that the solution becomes unattainable when the time step is too large. In fact, numerical noise typically accumulates during the integration of the ODEs [25], hence it is “non-local” in time. This is quite distinct from the physical noise in Eq. (9) which is instead local in time.
Fig. 1 illustrates how the percentage of solved instances and the EAMLLE vary with integration time step in problems with clause-to-variable ratio $M/N=7$. In these simulations, the system’s dynamics continued until a solution was reached or the simulation hit the maximum allowed number of steps. As expected, one observes that with an increase in the integration time step, the proportion of solved instances drops from 100% to 0%. The position of this drop correlates with the difficulty of the problem (characterized by the number of variables $N$): the solvable-unsolvable transition becomes sharper as the number of variables, $N$, increases. At this stage, however, we cannot conclude that such a transition is continuous in the “thermodynamic limit” ($N\rightarrow\infty$) or will become discontinuous (first order), since this distinction would require a theoretical analysis which is beyond the scope of this paper. However, we note that similar trends were observed in a recent paper [8], and in Sec. 3.3 we will show a similar phenomenon in the presence of physical noise.
According to our simulations, for the smallest $\Delta t$, the EAMLLE is negative across most instances. Fig. 1 shows that the EAMLLE increases monotonically with $\Delta t$. There is an interesting region, $0.05\lesssim\Delta t\lesssim 0.2$, where the positive EAMLLE co-exists with the 100 % solvability. We associate this region with a “transiently chaotic” dynamics regime. A shows LLE examples for two different time-step values.
@@
Figure 2: Power spectra for a satisfiable instance with 1000 variables and with clause-to-variable ratio $M/N=7$ generated using the integration time step of (a) 0.07, (b) 0.27, (c) 0.3, and (d) 0.4.
Fig. 2 shows an example of the power spectra for a typical instance. For $\Delta t=0.07$ we observe a prominent peak at lower frequencies (Fig. 2(a)), which gradually diminishes at larger time steps. We defer to Sec. 3.4 for the explanation of the origin of this peak since it does not involve noise (whether numerical or physical). Here, we just focus on the numerical noise effects. As the time step increases, the system ceases to reach a solution and becomes trapped in a possibly quasi-periodic orbit according to Fig. 2(d). In the Supplementary Information (SI), we provide power spectra for two other instances of the same size, see Figs. S1 and S2.
@@
Figure 3: Power spectra for a satisfiable instance with 90 variables and with clause-to-variable ratio $M/N=4.3$ obtained using the integration time step of (a) 0.05, (b) 0.25, (c) 0.54, and (d) 0.58.
@@
Similar results are obtained for other types of instances. For example, in Fig. 3 we present typical power spectra for instances with clause-to-variable ratio $M/N=4.3$, which are considered to be much harder to solve compared to $M/N=7$ instances [7]. A similar peak at low frequencies is observed for small $\Delta t$, which decreases as $\Delta t$ increases. The primary distinction for large instances is that, when $\Delta t$ is sufficiently large, the instance remains unsolved yet does not stall at the boundaries, there is a significant peak at zero frequency and a nearly flat high-frequency band that gradually increases, as shown on the inset of Fig. S1(d). Compared to Fig. 2, the curves in Fig. 3 are also noisier confirming that the instances in Fig. 3 are harder.
3.2 Physical noise
@@
We now move on to discuss the effect of physical noise as introduced according to the procedure discussed in Sec. 2.3. We then initialize all fast variables $v_{n}$’s to zero and increase the noise strength $\Gamma$. This is illustrated in Fig. 4, where the percentage of solved cases for a single problem instance (with $N=1000$ variables) with respect to the noise strength is plotted for clause-to-variable ratio $M/N=7$, along with the power spectrum of the fast variables’ flow vector field $\textnormal{d}\mathbf{v}/\textnormal{d}t$ for representative instances. For each value of $\Gamma$, 100 realizations of the noise are averaged over. A problem is considered unsolved if it does not reach a solution within 10,000 steps. In order to isolate the effect of noise, all instances are initialized identically. This choice is justified in Fig. S6, which demonstrates that using random voltages initial conditions and instances results in the same power spectrum behavior as using fixed ones. To minimize numerical errors, following the results presented in Fig. 1, we have chosen an integration time step $\Delta t$ of 0.05. (However, in Sec. 3.3, we consider an adaptive time step to further control the numerical errors while increasing the problem size.) The main plot shows the sharp decline in the success rate as noise increases, indicating the solvable-to-unsolvable transition.
@@
Figure 4: Main panel: Transition of the percentage of solved cases as a function of noise amplitude $\Gamma$, for a problem instance of 1000 variables at clause-to-variable ratio $M/N=7$. Each point represents 100 realization of the noise. In all cases, the variables $v_{n}$’s are initialized to zero, the long term memory $x_{l,m}$ to 1 and the short term memory $x_{s,m}$ to $C_{m}$ which is equal to 1/2 when $v_{n}=0$ $\forall n$ (midpoint value of the short term memory bounded range [0,1]). The spectrum is then averaged over the hundred noise realizations. The MLLE averaged over noise realisation $\langle\overline{\lambda}\rangle_{noise}$ is also displayed. Insets (1), (2), and (3) show representative power spectra of the flow vector field of the fast degrees of freedom of individual instances at different noise levels. Inset (1): Power spectrum of a solved instance at low noise ($\Gamma\approx 0$), where 100% of instances are solved. Inset (2): Power spectra at intermediate noise. The red curve corresponds to the average power spectrum of 85 noise realizations of solved cases, and the black curve to the average of 15 noise realizations of unsolved ones. Inset (3): Average power spectrum of an unsolved instance at high noise over 100 noise realizations ($\Gamma\approx 1.7$ where the solving rate drops to zero).
@@
As for the case of numerical noise, we also expect in this case that the power spectrum of the vector signal $\textnormal{d}\textbf{v}/\textnormal{d}t$ should have large discrete peaks at particular frequencies [37]. However, the spacing between neighboring frequency components should become negligible for aperiodic sequences like chaotic signals [38, 39]. The power spectrum of a chaotic solution should then contain all frequency components and to be broadband continuous without distinct periodicity [40]. This results in a spectrum that spreads continuously across frequencies and does not have any spikes.
@@
In agreement with the results obtained with numerical noise (Fig. 1), Fig 4 shows that the physical noise induces a decay in the percentage of solved cases. We again interpret this decay as the appearance of noise-induced chaos. Inset (1) of Fig 4 shows the power spectrum of the flow vector field $\textnormal{d}\mathbf{v}/\textnormal{d}t$ with no noise. There is a clear peak at low frequencies, showing a dominant frequency of the signal associated to regular dynamics. Inset (2) in red shows the power spectrum for the solved cases with intermediate noise strength. The peak is still there, but a drop at low frequencies due to $1/f$ noise appears. (We refer to B for a discussion of this type of noise.) The spectrum is also smoother than in Inset (1) because an average has been computed over all noise realizations. The sub-harmonics have canceled themselves out, and only the main frequency component remains. For the unsolved cases in black, $1/f$ noise is still present for small frequencies, and for larger frequencies there is a continuum spectrum, indicating the presence of chaos. The same features are found in the Inset (3), when no cases are solved. The amplitude of the power spectrum increases again for large frequencies due to the long term memory.
@@
Due to physical noise, the Lyapunov exponents are calculated using $v_{n}$’s that are not limited by fixed bounds as in Eq. (1). This is necessary because with strong noise, both the original state and the small perturbation could get stuck at the edges of the system, making it impossible to measure their separation. However, to keep the system stable, the norm of $\textnormal{d}\mathbf{v}/\textnormal{d}t$ is limited to a maximum of 10,000, so the $v_{n}$’s do not grow indefinitely. The quantity $\langle\overline{\lambda}\rangle_{noise}$ is obtained by averaging the largest Lyapunov exponents at each time step, and then averaging again over different noise realizations. This quantity is plotted in the main panel of Fig. 4 and its trend is very similar to the one we already discussed for numerical noise (cf. Fig. 1).
3.3 Transition as a function of problem size
@@
To study the transition into chaos with respect to the system size, an adaptive time step had to be used to better control numerical errors. The inverse of the largest time derivative of v is taken to scale the time step adaptively. This means that if the time derivative is large, the time step becomes small to prevent instability. Conversely, a smaller derivative leads to larger updates for faster progress. Therefore, $\Delta t$ is rescaled at every time step by a factor $1/(\text{max}\rvert{\text{d}\textbf{v}}/{\text{d}t}\lvert+\bar{\varepsilon})$, with $\bar{\varepsilon}=10^{-6}$ to avoid divergences when $\text{max}\rvert{\text{d}\textbf{v}}/{\text{d}t}\lvert=0$.
@@
The results are shown in Fig. 5. As expected, for low noise, all instances are solved. By increasing the noise strength, the system goes through a sort of “phase transition” where the number of successfully solved instances decreases dramatically. As in the case of numerical noise (cf. Fig. 1) this solvable-unsolvable transition becomes sharper as the number of variables, $N$, increases. [8] However, also in this case, it is not obvious if such a transition will remain continuous in the “thermodynamic limit” ($N\rightarrow\infty$) or will become discontinuous (first order). Such a theoretical analysis is beyond the scope of this paper.
@@
Figure 5: Percentage of problem instances solved with respect to the noise strength $\Gamma$ for clause-to-variable ratio $M/N=7$. The solvable-unsolvable transition becomes sharper with increasing number of variables. The time step is set to be inversely proportional to the maximum time derivative of the variables’ vector v (see main text for details). For every value of $\Gamma$ and for each size $N$ a batch of 1000 problem instances has been simulated in one run.
3.4 Explanation of the peaks in the spectra
(a) Refer to caption(b) Refer to caption
@@ 3.4 Explanation of the peaks in the spectra
Figure 6: Evolution of the averaged power spectra for different parameter settings in simulations with no noise. The spectra are averaged over 10 different problem instances of $N=1000$, clause-to-variables ratio of $M/N=7$, and time step $\Delta t$ of 0.05. (a): Variation of $\alpha$ with fixed $\beta=20$. As $\alpha$ increases, the spectrum peak amplitude grows indicating stronger oscillatory behavior, while the peak frequency remains approximately constant (b): Variation of $\beta$ with fixed $\alpha=5$. The spectrum peak shifts to higher frequencies and its amplitude grows with increasing $\beta$, revealing the system’s strong sensitivity to this parameter in controlling oscillation frequency.
@@
Finally, to understand the origin of the peaks in the power spectra, we refer to Eqs. (2) and (3), in the absence of both physical and numerical noise. These equations show that $\beta$ and $\alpha$ control the rates of change of the memory variables $x_{s,m}$ and $x_{l,m}$, respectively. We report in Fig. 6 simulations performed without physical noise and with all other parameters held constant. The power spectra are averaged over 10 different problem instances. As shown in Fig. 6, increasing $\beta$ causes the peak of the power spectrum to shift noticeably toward higher frequencies, whereas increasing $\alpha$ appears to have little to no effect on its position. The difference is that $x_{s,m}$ appears in its own rate equation, creating a self-reinforcing feedback loop, while $x_{l,m}$ does not appear in its own rate equation. Thus, $\beta$ controls a state-dependent feedback mechanism in Eq. (2) inducing faster fluctuations in the variables’ Eq. (1), where $\beta$ acts indirectly through multiplicative influence on $x_{s,m}$. In contrast, increasing $\alpha$ has little effect on the spectral peak: while it accelerates the dynamics of $x_{l,m}$ in Eq. (3), this term is not multiplied by any internal state variable and contributes more smoothly to the varaiables $v_{n}$’s dynamics. Consequently, its influence does not substantially shift the dominant frequencies in $\dot{v}_{n}$.
4 Discussion

### v2

@@ II.3 Physical noise
To make contact with experimental realizations of these machines, we introduce physical noise through stochastic terms in the ODEs of DMMs. In view of the fact that in possible electrical circuit realizations of DMMs, the memory variables are the ones most susceptible to noise Traversa and Di Ventra (2017a), we add stochastic terms to Eqs. (2) and (3). Of course, such noise propagates indirectly to the fast degrees of freedom $\textnormal{v}_{n}$’s through their coupling in Eq. (1). Indeed, as discussed in Bearden et al. (2018), noise can, e.g., serve as a proxy for temperature, since temperature fluctuations can induce physical instabilities in the resistive memory elements of the gates of DMMs.
@@
If we indicate with $\tilde{y}_{m}$ the noisy version of the long-term memory variables, the stochastic terms are linked to the noiseless dynamical system’s memory by
$\dot{\tilde{y}}_{m}(t)=\dot{y}_{m}(t)+\Gamma\eta(t)$ (9)
@@
with $\eta\sim\mathcal{N}(0,1)$, a Gaussian distribution with zero average and unit variance, and $\Gamma$ is the strength of the noise. Similarly for the short term memory $x_{m}$.
@@
Figure 1: Numerical noise investigation: Percentage of solved instances and EAMLLE ($\langle\overline{\lambda}\rangle$) vs. integration time step, $\Delta t$. Each point used to make this plot was obtained using an ensemble of 100 solvable 3-SAT instances with clause-to-variable ratio $M/N=7$. The upper bound for each trajectory was set to 20,000 steps. The smallest $\Delta t$ taken is 0.01.
III Results
III.1 Numerical noise
@@
By numerically integrating Eqs. (1)-(3) using the forward Euler method, we transform the continuous-time dynamics into a discrete map. We view related errors induced by such a transformation as the primary components of numerical noise, which can sometimes force the system into chaotic behavior. It is clear that the solution becomes unattainable when the time step is too large. In fact, numerical noise typically accumulates during the integration of the ODEs Press et al. (2007), hence it is “non-local” in time. This is quite distinct from the physical noise in Eq. (9) which is instead local in time.
@@
Figure 2: Numerical noise investigation: Power spectra for a satisfiable instance with $N=1000$ variables and with clause-to-variable ratio $M/N=7$ generated using the integration time step of (a) 0.07, (b) 0.2, (c) 0.3, and (d) 0.4.
Fig. 1 illustrates how the percentage of solved instances and the EAMLLE vary with integration time step in problems with clause-to-variable ratio $M/N=7$. In these simulations, the system’s dynamics continued until a solution was reached or the simulation hit the maximum allowed number of steps. We note that the total time of integration varies according to the instances to allow enough time to reach a solution of the 3-SAT problem. As expected, one observes that with an increase in the integration time step, the proportion of solved instances drops from 100% to 0%. The position of this drop correlates with the difficulty of the problem (characterized by the number of variables $N$): the solvable-unsolvable transition becomes sharper as the number of variables, $N$, increases. At this stage, however, we cannot conclude that such a transition is continuous in the “thermodynamic limit” ($N\rightarrow\infty$) or will become discontinuous (first order), since this distinction would require a theoretical analysis which is beyond the scope of this paper. However, we note that similar trends were observed in a recent paper Zhang and Di Ventra (2021), and in Sec. III.3 we will show a similar phenomenon in the presence of physical noise.
According to our simulations, for the smallest $\Delta t$, the EAMLLE is negative across most instances. Fig. 1 shows that the EAMLLE increases monotonically with $\Delta t$. However, there is an interesting region, $0.05\lesssim\Delta t\lesssim 0.2$, where the positive EAMLLE co-exists with the 100 % solvability. We associate this region with a “transiently chaotic” dynamics regime. Appendix B shows LLE examples for two different time-step values.
@@
Figure 3: Numerical noise investigation: Power spectra for a satisfiable instance with $N=90$ variables and with clause-to-variable ratio $M/N=4.3$ obtained using the integration time step of (a) 0.05, (b) 0.25, (c) 0.54, and (d) 0.58.
Fig. 2 shows an example of the power spectra for a typical instance. For $\Delta t=0.07$ we observe a prominent peak at lower frequencies (Fig. 2(a)), which gradually diminishes at larger time steps. We defer to Sec. III.4 for the explanation of the origin of this peak since it does not involve noise (whether numerical or physical). Here, we just focus on the numerical noise effects. As the time step increases, the system ceases to reach a solution and becomes trapped in a possibly quasi-periodic orbit according to Fig. 2(d). In the Supplementary Information (SI), we provide power spectra for two other instances of the same size, see Figs. S1 and S2.
@@
Similar results are obtained for other types of instances. For example, in Fig. 3 we present typical power spectra for instances with clause-to-variable ratio $M/N=4.3$, which are considered to be much harder to solve compared to $M/N=7$ instances Bearden, Pei, and Di Ventra (2020). A similar peak at low frequencies is observed for small $\Delta t$, which decreases as $\Delta t$ increases. The primary distinction for large instances is that, when $\Delta t$ is sufficiently large, the instance remains unsolved yet does not stall at the boundaries, there is a significant peak at zero frequency and a nearly flat high-frequency band that gradually increases, as shown on the inset of Fig. S1(d). Compared to Fig. 2, the curves in Fig. 3 are also noisier confirming that the instances are harder.
III.2 Physical noise
@@
We now move on to discuss the effect of physical noise as introduced according to the procedure discussed in Sec. II.3. We then initialize all fast variables $\textnormal{v}_{n}$’s to zero and increase the noise strength $\Gamma$. This is illustrated in Fig. 4, where the percentage of solved cases for a single problem instance (with $N=1000$ variables) with respect to the noise strength is plotted for clause-to-variable ratio $M/N=7$, along with the power spectrum of the fast variables’ flow vector field $\textnormal{d}\mathbf{v}/\textnormal{d}t$ for representative instances. A problem is considered unsolved if it does not reach a solution within $10^{4}$ steps. The physical noise introduced in Eq. (9) is genuinely stochastic, so each trajectory corresponds to a distinct realization of the random process $\eta(t)$, while the problem instance and initial conditions are kept fixed. Ensemble-averaged quantities such as $\langle\overline{\lambda}\rangle_{\text{noise}}$ are therefore obtained by averaging over 100 independent noise realizations. In contrast, the numerical noise in Sec. III A is deterministic once $\Delta t$ is chosen, so there is no concept of “realizations,” and averaging is instead performed over different problem instances. The choice of initializing all the instances identically is justified in Fig. S6, which demonstrates that using random initial conditions and instances results in the same power spectrum behavior as using fixed ones. To minimize numerical errors, following the results presented in Fig. 1, we have chosen an integration time step $\Delta t$ of 0.05. (However, in Sec. III.3, we consider an adaptive time step to further control the numerical errors while increasing the problem size.) The main plot of Fig. 4 shows the sharp decline in the success rate as noise increases, indicating the solvable-to-unsolvable transition.
@@
Figure 4: Physical noise investigation. Main panel: Transition of the percentage of solved cases as a function of noise amplitude $\Gamma$, for a problem instance of $N=1000$ variables at clause-to-variable ratio $M/N=7$. Each point represents 100 realization of the noise. The simulation runs for a maximum of 10,000 steps if convergence to a solution does not occur beforehand . In all cases, the variables $\textnormal{v}_{n}$’s are initialized to zero, the long term memory $y_{m}$ to 1 and the short term memory $x_{m}$ to $C_{m}$ which is equal to 1/2 when $\textnormal{v}_{n}=0$ $\forall n$ (midpoint value of the short term memory bounded range [0,1]). The spectrum is then averaged over 100 noise realizations. The MLLE averaged over noise realisation $\langle\overline{\lambda}\rangle_{noise}$ is also displayed. Insets (1), (2), and (3) show representative power spectra of the flow vector field of the fast degrees of freedom of individual instances at different noise levels. Inset (1): Power spectrum of a solved instance at low noise ($\Gamma\approx 0$), where 100% of instances are solved. Inset (2): Power spectra at intermediate noise. The red curve corresponds to the average power spectrum of noise realizations of 85 solved cases, and the black curve to the average of noise realizations of 15 unsolved ones. Inset (3): Average power spectrum of an unsolved instance at high noise over 100 noise realizations ($\Gamma\approx 1.7$ where the solving rate drops to zero).
@@
As for the case of numerical noise, we also expect in this case that the power spectrum of the vector signal $\textnormal{d}\textbf{v}/\textnormal{d}t$ should have large discrete peaks at particular frequencies Melbourne and Gottwald (2008). However, the spacing between neighboring frequency components should become negligible for aperiodic sequences like chaotic signals H. G. Schuster (2005); Rasband (1990). The power spectrum of a chaotic solution should then contain all frequency components and to be broadband continuous without distinct periodicity Lipton and Dabke (1996). This results in a spectrum that spreads continuously across frequencies and does not have any spikes.
@@
In agreement with the results obtained with numerical noise (Fig. 1), Fig 4 shows that the physical noise induces a decay in the percentage of solved cases. We again interpret this decay as the appearance of noise-induced chaos. Inset (1) of Fig 4 shows the power spectrum of the flow vector field $\textnormal{d}\mathbf{v}/\textnormal{d}t$ with no noise. There is a clear peak at low frequencies, showing a dominant frequency of the signal associated to regular dynamics. Inset (2) in red shows the power spectrum for the solved cases with intermediate noise strength. The peak is still there, but a drop at low frequencies due to $1/f$ noise appears. (We refer to Appendix  C for a discussion of this type of noise.) The spectrum is also smoother than in Inset (1) because an average has been computed over all noise realizations. The sub-harmonics have canceled themselves out, and only the main frequency component remains. For the unsolved cases in black, $1/f$ noise is still present for small frequencies, and for larger frequencies there is a continuum spectrum, indicating the presence of chaos. The same features are found in the Inset (3), when no cases are solved. The amplitude of the power spectrum increases again for large frequencies due to the long term memory.
@@
Due to physical noise, the Lyapunov exponents are calculated using $\textnormal{v}_{n}$’s that are not limited by fixed bounds as in Eq. (1). This is necessary because with strong noise, both the original state and the small perturbation could get stuck at the edges of the system, making it impossible to measure their separation. However, to keep the system stable, the norm of $\textnormal{d}\mathbf{v}/\textnormal{d}t$ is limited to a maximum of $10^{4}$, so the $\textnormal{v}_{n}$’s do not grow indefinitely. The quantity $\langle\overline{\lambda}\rangle_{noise}$ is obtained by averaging the largest Lyapunov exponents at each time step, and then averaging again over different noise realizations. This quantity is plotted in the main panel of Fig. 4 and its trend is very similar to the one we already discussed for numerical noise (cf. Fig. 1). We emphasize again that the reported Lyapunov exponents correspond to finite-time, while the asymptotic value remains zero below a critical noise strength. According to the standard picture of noise-induced chaos, small stochastic perturbations act to sustain the transient chaos already present in the noise-free system. The noise continually kicks trajectories away from the regular attractor, allowing them to revisit the unstable region near the chaotic saddle. As the noise amplitude increases, trajectories spend progressively longer time near this saddle; beyond a critical noise strength, the random kicks prevent escape altogether, effectively transforming the chaotic saddle into a true chaotic attractor. This mechanism naturally accounts for the observed growth of finite-time Lyapunov exponents with increasing noise strength in our simulations.
III.3 Transition to chaos as a function of problem size
@@
To study the transition into chaos with respect to the system size, an adaptive time step had to be used to better control numerical errors. The inverse of the largest time derivative of v is taken to scale the time step adaptively. This means that if the time derivative is large, the time step becomes small to prevent instability. Conversely, a smaller derivative leads to larger updates for faster progress. Therefore, $\Delta t$ is rescaled at every time step by a factor of $1/(\text{max}\rvert{\text{d}\textbf{v}}/{\text{d}t}\lvert+\bar{\varepsilon})$, with $\bar{\varepsilon}=10^{-6}$ to avoid divergences when $\text{max}\rvert{\text{d}\textbf{v}}/{\text{d}t}\lvert=0$.
@@
The results are shown in Fig. 5. As expected, for low noise, all instances are solved. By increasing the noise strength, the system goes through a sort of “phase transition” where the number of successfully solved instances decreases dramatically. As in the case of numerical noise (cf. Fig. 1) this solvable-unsolvable transition becomes sharper as the number of variables, $N$, increases. Zhang and Di Ventra (2021) However, also in this case, it is not obvious if such a transition will remain continuous in the “thermodynamic limit” ($N\rightarrow\infty$) or will become discontinuous (first order). Such a theoretical analysis is beyond the scope of this paper.
@@
Figure 5: Physical noise investigation: Percentage of problem instances solved with respect to the noise strength $\Gamma$ for clause-to-variable ratio $M/N=7$. The solvable-unsolvable transition becomes sharper with increasing number of variables. The time step is set to be inversely proportional to the maximum time derivative of the variables’ vector v (see main text for details). For every value of $\Gamma$ and for each size NN a batch of 1000 problem instances has been simulated in one run.
III.4 Explanation of the peaks in the spectra
(a) Refer to caption
(b) Refer to caption
@@ III.4 Explanation of the peaks in the spectra
Figure 6: Evolution of the averaged power spectra for different parameter settings in simulations with no noise. The spectra are averaged over 10 different problem instances of $N=1000$, clause-to-variables ratio of $M/N=7$, and time step $\Delta t$ of 0.05. (a): Variation of $\alpha$ with fixed $\beta=25$. As $\alpha$ increases, the spectrum peak amplitude grows indicating stronger oscillatory behavior, while the peak frequency remains approximately constant (b): Variation of $\beta$ with fixed $\alpha=5$. The spectrum peak shifts to higher frequencies and its amplitude grows with increasing $\beta$, revealing the system’s strong sensitivity to this parameter in controlling oscillation frequency.
@@
Finally, to understand the origin of the peaks in the power spectra, we refer to Eqs. (2) and (3), in the absence of both physical and numerical noise. These equations show that $\beta$ and $\alpha$ control the rates of change of the memory variables $x_{m}$ and $y_{m}$, respectively. We report in Fig. 6 simulations performed without physical noise and with all other parameters held constant. The power spectra are averaged over 10 different problem instances. As shown in Fig. 6, increasing $\beta$ causes the peak of the power spectrum to shift noticeably toward higher frequencies, whereas increasing $\alpha$ appears to have little to no effect on its position. The difference is that $x_{m}$ appears in its own rate equation, creating a self-reinforcing feedback loop, while $y_{m}$ does not appear in its own rate equation. Thus, $\beta$ controls a state-dependent feedback mechanism in Eq. (2) inducing faster fluctuations in the variables of Eq. (1), where $\beta$ acts indirectly through multiplicative influence on $x_{m}$. In contrast, increasing $\alpha$ has little effect on the spectral peak: while it accelerates the dynamics of $y_{m}$ in Eq. (3), this term is not multiplied by any internal state variable and contributes more smoothly to the varaiables $\textnormal{v}_{n}$’s dynamics. Consequently, its influence does not substantially shift the dominant frequencies in $\dot{\textnormal{v}}_{n}$.
IV Discussion